OpenStudy (anonymous):

show sin^2x(1+sin^2x)=sec^2x-cos^2x

OpenStudy (cwrw238):

try substituting sin^2x = 1 - cos^2x in left hand side

OpenStudy (anonymous):

then I get:\[1+\sec ^{2}\theta -\cos ^{2}\theta-\sec ^{2}\theta \cos ^{2}\theta \]

OpenStudy (cwrw238):

right - now replace cos^2x by 1 /sec^2x in the last term and what do u get?

OpenStudy (anonymous):

OMG! Thankss!

OpenStudy (cwrw238):

lol - welcome

OpenStudy (anonymous):

But when I opened up the brackets at the start it didn't work out, can you check why?

OpenStudy (cwrw238):

lol - to be perfectly honest i went wrong myself thats why i abanded it - i'll have another look how did u get your result?

OpenStudy (anonymous):

I just followed your way, 1st sub. sin^2x = 1-cos^2x and then in the last term subing cos^2x by 1/sec^2x and then cancelling it outt

OpenStudy (cwrw238):

it can't be a good day for me because i can't get 1+sec^2θ−cos^2θ−sec^2θcos^2θ

OpenStudy (cwrw238):

i'll have to come back to u on this...

OpenStudy (cwrw238):

@sim123 try plugging in x = 30 degrees into each side of the identity and see what you get. i'll try it as well.

OpenStudy (cwrw238):

sin^2 30 (1 + sin^2 30) = 0.5^2(1 + 0.5^2) = 0.25 * 1.25 = 0.3125

OpenStudy (cwrw238):

sec^2 30 - cos^2 30 = 1.33333 - 0.75 = 0.583333

OpenStudy (cwrw238):

we've been hitting our heads against a brick wall - they are not equal